OpenStudy (anonymous):

Let (theta) be an angle in quadrant 2 such that cos(theta) = -4/5 . Find the exact values of csc(theta) and tan(theta)

OpenStudy (campbell_st):

just draw a triangle in the quadrant  then you can find csc and tan

OpenStudy (anonymous):

Awesome, almost had this but posted it just incase. I got \[\csc \theta\] = 5/3 and \[\tan \theta\] = -3/4

OpenStudy (campbell_st):

yep... thats it